# Explainable Differential Privacy-Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

### v3

@@ Abstract
Machine Learning (ML) models combined with in-situ sensing offer a powerful solution to address defect detection challenges in Additive Manufacturing (AM), yet this integration raises critical data privacy concerns, such as data leakage and sensor data compromise, potentially exposing sensitive information about part design and material composition. Differential Privacy (DP), which adds mathematically controlled noise to ML models, provides a way to balance data utility with privacy by concealing identifiable traces from sensor data. However, introducing noise into ML models, especially black-box Artificial Intelligence (AI) models, complicates the prediction of how noise impacts model accuracy. This study presents the Differential Privacy-Hyperdimensional Computing (DP-HD) framework, which leverages Explainable AI (XAI) and the vector symbolic paradigm to quantify noise effects on accuracy. By defining a Signal-to-Noise Ratio (SNR) metric, DP-HD assesses the contribution of training data relative to DP noise, allowing selection of an optimal balance between accuracy and privacy. Experimental results using high-speed melt pool data for anomaly detection in AM demonstrate that DP-HD achieves superior operational efficiency, prediction accuracy, and privacy protection. For instance, with a privacy budget set at 1, DP-HD achieves $94.43\%$ accuracy, outperforming state-of-the-art ML models. Furthermore, DP-HD maintains high accuracy under substantial noise additions to enhance privacy, unlike current models that experience significant accuracy declines under stringent privacy constraints.
@@ 1 Introduction
To effectively anticipate the impact of DP noise on model accuracy, this paper introduces the Differential Privacy-Hyperdimensional Computing (DP-HD) framework. This explainable approach is tailored for in-process monitoring, featuring rapid query response capabilities, making it ideal for applications requiring swift analysis while preserving both privacy and performance fidelity. The strength of DP-HD lies in its innovative data processing method, which utilizes high-dimensional vector spaces for computation. This design not only enables efficient analysis but also provides a way to define privacy levels precisely and predict model accuracy before adding noise, aided by the introduction of a Signal-to-Noise Ratio (SNR) metric to measure the balance between data contribution and DP noise. The primary contributions of this work are as follows:
@@
    We propose an explainable DP-HD framework to improve transparency in interactions between ML models and DP. The framework introduces a method for determining the optimal Standard Deviation (SD) for generating random basis vectors, aided by the use of an SNR metric to achieve an ideal balance between privacy protection and performance accuracy.
@@
    We assess the impact of noise on the framework’s memorization capacity, utilizing SNR to showcase DP-HD’s ability to predict noise effects on model accuracy across various privacy levels.
@@
    We present an inference privacy technique that mitigates the risk of decoding query hypervectors by removing low-variance dimensions, further enhancing the security of DP-HD during inference.
@@ 2.2 Privacy-Preserving in In-situ Sensing
The state-of-the-art method for applying DP to ML models is Differentially Private Stochastic Gradient Descent (DP-SGD), a variation of the standard Stochastic Gradient Descent (SGD) algorithm. DP-SGD enhances privacy by introducing precisely calibrated noise to the gradients during the training process. This technique minimizes the impact of individual data entries by clipping gradients and adding noise, thereby masking their specific contributions to the final model parameters. DP-SGD provides measurable privacy assurances, essential in fields where data confidentiality is paramount [62].
2.3 Explainable AI
@@ 3.1 Hyperdimensional Computing
Hyperdimensional Computing (HD) draws significant inspiration from the architecture of the human brain, utilizing high-dimensional vectors to simulate cognitive functions [76]. This advanced approach allows HD to emulate brain-like methodologies for processing, analyzing, and storing information across diverse cognitive tasks. The HD framework, depicted in Figure 1, is structured into four key phases: encoding, training, inference, and retraining. Each phase is meticulously designed to model different aspects of cognitive processing.
@@ 3.3 Differential Privacy-Hyperdimensional Computing
In this work, we present the DP-HD model, a novel approach in AM that incorporates DP. This model excels in real-time monitoring and allows for precise mathematical adjustments of noise levels, essential for safeguarding privacy without compromising accuracy, as illustrated in Algorithm 1. Ensuring the confidentiality of training data in AM is crucial to prevent potential security breaches. To address this, Gaussian noise is strategically injected into the class hypervectors, thereby preventing the extraction of individual training samples by unauthorized entities.
During the training phase of the HD model, DP is integrated by making the class hypervectors differentially private through the addition of Gaussian noise, as depicted in Equation (8). This approach effectively reduces the likelihood of an adversary inferring specific details about the original feature vectors from the class hypervectors. The optimal $\sigma_{dp}$ is determined using Equation (7), which aims to balance data privacy and the operational performance of our predictive model in the AM setting. In our specific case, we set $\delta$ to $10^{-4}$, based on the principle that $\delta$ should be smaller than the inverse of the dataset size. The selection of the appropriate noise level, which is crucial for enhancing the model’s security, depends on the DP-HD model’s sensitivity. This sensitivity is determined by identifying the maximum norm across all encoded training samples, as shown in Equation (9). A higher sensitivity necessitates the incorporation of greater noise levels, thus reinforcing the model’s security. Therefore, the highest sensitivity observed across all training samples is used as a collective sensitivity indicator, ensuring robust privacy measures across the dataset.
@@
Another crucial parameter that dictates the privacy level and sets the noise level is the privacy budget. The privacy budget can be selected to predict the impact of noise on the model’s performance. In Section 3.3.1, we delve into the explainability of the DP-HD mechanism. We begin by assessing the effect of noise and training samples on the model’s performance through an examination of its memorization capability. We then define the SNR, which is derived from the ratio of the signal (training samples) contribution to the model’s performance against the noise contribution. This metric enables us to forecast how noise levels influence accuracy, allowing us to set specific privacy and performance levels concurrently. Following this, we explore the decodability of DP-HD. We analyze an analytical method for extracting information from a hypervector by subtracting two secure class hypervectors generated from datasets differing by only one sample. Moreover, in Section 3.3.2, we outline an algorithm to find the optimal encoding strategy to balance privacy and performance. The encoding phase is pivotal to the model’s success, as it determines how data is mapped from feature space to hyperspace.
@@ 3.3.1 Explainability
Memorization in DP-HD occurs through the bundling operation during the training phase. As described earlier, the memorized vector is given by Equation (1). To ensure DP during training, Gaussian noise with an SD of $\Delta g\sigma_{dp}$ is added to each component of the class hypervectors, as shown in Equation (8). This noise plays a crucial role during the inference process, where the algorithm attempts to recall one of the memorized hypervectors. In the presence of noise, the inference primarily depends on the similarity inner product between the secured class hypervector $\tilde{\vec{C}}_{s}$ and the query hypervector $\vec{H_{q}}$:
@@
This relationship allows us to quantify the impact of noise on the system’s accuracy. A higher SNR indicates better accuracy by emphasizing the signal’s dominance over noise during the inference stage. This equation also shows that increasing $\epsilon$ improves the SNR, resulting in higher accuracy. However, it also reduces the noise level, thereby decreasing the privacy level according to Equations (7) and (8). Therefore, we select an $\epsilon$ value that balances the required privacy and model performance levels.
Moreover, as we can see, with a very large dimension $D$, the SNR ratio drastically decreases. The only way to counter this effect is to increase $\sigma_{dp}$, which is effectively done through a loss of DP by increasing $\epsilon$ and $\delta$. However, an alternate method of increasing the SNR is by simply increasing the dataset size $N$. Thus, having more training data reduces the effect of DP on the quality of the model. This can be intuitively understood since the data is sampled from the same underlying data distribution, the more dense the training sample is, the less the effect of removing a single training point becomes. In the limit of small ${\rm SNR}$, for good performance the width of the signal distribution should be much smaller than the width of the noise distribution.
@@
$=\mathbb{E}\left((A_{i})^{-1}\vec{\xi}_{i}\vec{\xi}_{i}^{\dagger}(A_{i}^{\dagger})^{-1}\right)$ (20)
$=\sqrt{2}\mathbb{E}\left((A_{i})^{-1}\vec{\xi}_{i}\vec{\xi}_{i}^{\dagger}(A_{i}^{\dagger})^{-1}\right)$ (21)
@@
$\hat{\vec{F}}=\vec{F}+N(\vec{0},\Sigma)$ (23)
where $\Sigma=\Sigma_{0}/K={\sqrt{2}\Delta g\sigma_{dp}}/K$. Thus, as expected, the noise during the decoding process also depends linearly on the sensitivity of the encoder and the amount of differential private noise added to the encoder. If we want better DP, then we need to add more noise which will increase the error in the decoder. This error can be reduced by increasing the value of the sampled in $K$. However, note that since $\Delta g$ is proportional to $D$, and $K$ is proportional to $D$, the ratio $\Delta g/K$ is independent of $D$. Thus, increasing the dimension should not change the error in the decoding process.
@@ 3.3.2 Encoding Strategy
The encoding strategy plays a vital role in balancing privacy and performance in the DP-HD model. To achieve this balance, it is essential to find the optimal SD, denoted as ${\sigma}^{*}_{\text{b}}$, which ensures both the highest accuracy and the highest privacy. The value of ${\sigma}^{*}_{\text{b}}$ is influenced by the distribution of training samples and the distance between the centers of different classes. It is important to note that ${\sigma}^{*}_{\text{b}}$ is not dependent on the DP-HD model parameters, such as the size of the hypervectors.
To determine ${\sigma}^{*}_{\text{b}}$, we select a subset of training samples that reflect the distribution of the entire dataset for training, and another subset for validation. We then experiment with different ${\sigma}_{\text{b}}$ values to generate random basis vectors, as outlined in Algorithm 2. Using these values, we create secured class hypervectors with the training data for various privacy budget values. We then calculate the validation accuracy based on the model’s performance in predicting the validation data labels. We store the accuracy for each pair of (${\sigma}_{\text{b}}$,$\epsilon$). Finally, we select the SD for generating random basis vectors, ${\sigma}^{*}_{\text{b}}$, that corresponds to the highest accuracy and the lowest $\epsilon$.
@@ 3.4.1 Decodability of Hypervectors
Figure 3: Schematic view of the experiment, including (a) camera setup positioned in the build chamber’s upper right corner, (b) the test artifact in EOS M270 laser powder bed fusion AM machine, (c) top view of the part, and (d) side view dimensions of the part with 40.5-degree overhang.
@@ 4 Experimental Design
Addressing privacy concerns in the dissemination and analysis of sensor data from AM processes is a paramount consideration. The need to protect sensitive manufacturing data, especially in an era where intellectual property and competitive advantage are closely guarded, necessitates the exploration of privacy-preserving methodologies. This study, therefore, not only focuses on the technical aspects of in-situ sensing and defect detection but also realizes the implementation of DP measures. These measures ensure that the valuable insights gained from high-speed camera data do not compromise the confidentiality of the manufacturing process or the design specifications of the parts being produced. This introduction sets the stage for a comprehensive investigation into the experimental designs and methodologies employed in this research. The subsequent sections will delve into the specifics of sensor integration, data acquisition, and the analytical techniques used to achieve the dual objectives of enhancing manufacturing precision and ensuring data privacy in the realm of AM.
@@ 4.1 LPBF AM Machine Instrumentation and Setup
This section outlines the instrumentation of a commercial LPBF machine, specifically the EOS M270 model situated at the National Institute of Standards and Technology (NIST). Integration involved outfitting the machine with a high-speed visible camera. The camera boasts a frame rate of 400 frames per second and a spectral response spanning 300-950 nm, facilitating real-time monitoring of the LPBF process. This monitoring entails capturing disparities in the thermal signature of the part as it undergoes layer-wise construction, enhancing quality assurance. Positioned within the build chamber’s upper right corner, the high-speed visible camera is connected via data cables routed through a custom sealed through-port. Notably, precise measurements regarding the camera’s orientation and distance were not conducted. The camera utilized features a Silicon-based array and boasts a resolution of 1.2 megapixels. Its placement inside the build chamber enables comprehensive imaging of the part’s shape and surrounding spatter pattern. Images captured by the high-speed visible camera are windowed to dimensions of 256 pixels × 256 pixels and acquired at a rate of 1000 frames per second. These images serve as representative snapshots, depicting both overhang and bulk build features. The experimental focus lies on a straightforward unsupported overhang geometry, selected by NIST researchers for its suitability in investigating meltpool behavior during overhang construction, thereby facilitating subsequent physical modeling efforts [78].
@@
The example part was fabricated utilizing specific build parameters tailored for a commercial LPBF system. These parameters include a hatch distance of 0.1 mm, a stripe width of 4 mm, a stripe overlap of 0.1 mm, a powder layer thickness of 20 $\mu$m, a laser scan speed of 800 mm/s, and a laser power of 195 W during infill. Contour passes commence from distinct corners of the part, as indicated in Figure 3(c), traversing counterclockwise along the part perimeter. Laser power settings for these passes vary, with pre-contour passes utilizing 100 W and post-contour passes employing 120 W, both at a scan speed of 800 mm/s. The study’s focus lies on acquiring and initially processing sensor data, which entails exploring a single layer example from the 16 mm tall part. Data collection spans 45 different build layers, with the overhang delineated as the last two scan vectors prior to or just after edge formation, excluding pre- or post-contour scans. The remaining scans are attributed to the bulk volume of the part. A stripe pattern scan strategy is adopted (Figure 3(c) and 3(d)), with laser scans over the overhang occurring four times per layer beyond 4 mm build height. The stripe orientation shifts by 90° between layers, and three example layers exhibit a vertical stripe pattern. The test artifact, composed of nickel alloy 625, features a 40.5-degree overhang without additional support structures. Sensor data analysis focuses on three example build heights: 6.06 mm, 7.90 mm, and 9.70 mm, encompassing the formation of the overhang structure. Each layer consists of four 4 mm stripes alternated by 90°. The nickel alloy 625 powder originates from the machine vendor, with a mean powder diameter of 37.8 $\mu$m, utilized in conjunction with a custom alloy 625 substrate. A side-view schematic of the LPBF part (Figure 3(d)) illustrates the investigation into the effects of an overhanging structure. Data exploration in this paper centers on a single build layer at a height of 7.90 mm, although the overhanging structure extends to complete a 16 mm tall object.

### v4

@@ Abstract
Machine Learning (ML) models integrated with in-situ sensing offer transformative solutions for defect detection in Additive Manufacturing (AM), but this integration brings critical challenges in safeguarding sensitive data, such as part designs and material compositions. Differential Privacy (DP), which introduces mathematically controlled noise, provides a balance between data utility and privacy. However, black-box Artificial Intelligence (AI) models often obscure how this noise impacts model accuracy, complicating the optimization of privacy–accuracy trade-offs. This study introduces the Differential Privacy-Hyperdimensional Computing (DP-HD) framework, a novel approach combining Explainable AI (XAI) and vector symbolic paradigms to quantify and predict noise effects on accuracy using a Signal-to-Noise Ratio (SNR) metric. DP-HD enables precise tuning of DP noise levels, ensuring an optimal balance between privacy and performance. The framework has been validated using real-world AM data, demonstrating its applicability to industrial environments. Experimental results demonstrate DP-HD’s capability to achieve state-of the-art accuracy (94.43%) with robust privacy protections in anomaly detection for AM, even under significant noise conditions. Beyond AM, DP-HD holds substantial promise for broader applications in privacy-sensitive domains such as healthcare, financial services, and government data management, where securing sensitive data while maintaining high ML performance is paramount.
@@ 1 Introduction
To effectively anticipate the influence of DP noise on model accuracy, this paper introduces the Differential Privacy-Hyperdimensional Computing (DP-HD) framework. This explainable approach is tailored for in-process monitoring, featuring rapid query response capabilities, making it ideal for applications requiring swift evaluation while preserving both privacy and performance fidelity. The strength of DP-HD lies in its innovative data processing method, which utilizes high-dimensional vector spaces for computation. This design not only enables efficient analysis but also provides a way to define privacy levels precisely and predict model accuracy before adding noise, aided by the introduction of a Signal-to-Noise Ratio (SNR) metric to measure the balance between data contribution and DP noise. A case study using real-world AM data, representative of typical industrial conditions, demonstrates the practical benefits of DP-HD, showing its ability to safeguard sensitive data while maintaining high accuracy in defect detection. Beyond AM, this framework offers potential for other privacy-sensitive industries, such as aerospace and healthcare, where protecting proprietary data is critical. The primary contributions of this work are as follows:
@@
    We propose an explainable DP-HD framework to enhance transparency in interactions between ML models and DP. The framework introduces a method for determining the optimal Standard Deviation (SD) for generating random basis vectors, aided by the use of an SNR metric to achieve an ideal balance between privacy protection and performance accuracy.
@@
    We evaluate the impact of noise on the framework’s memorization capacity, demonstrating how the SNR metric enables DP-HD to accurately predict noise effects on model accuracy across various privacy levels, a critical capability for privacy-sensitive applications.
@@
    We present a novel inference privacy technique that reduces the risk of decoding query hypervectors by removing low-variance dimensions. This technique significantly enhances the security of DP-HD, making it robust and scalable for real-world industrial applications.
@@ 2.2 Privacy-Preserving in In-situ Sensing
The state-of-the-art approach for applying DP to ML models is Differentially Private Stochastic Gradient Descent (DP-SGD), a variation of the standard Stochastic Gradient Descent (SGD) algorithm. DP-SGD enhances privacy by introducing precisely calibrated noise to the gradients during the training process. This technique minimizes the impact of individual data entries by clipping gradients and adding noise, thereby obscuring their specific contributions to the final model parameters. DP-SGD provides measurable privacy assurances, essential in fields where data confidentiality is paramount [62].
@@
Another approach, Rényi Differential Privacy (RDP), extends the DP framework by quantifying privacy loss using Rényi divergence, which measures the statistical distance between probability distributions. RDP introduces a more flexible way to compose privacy guarantees over multiple queries by leveraging the additivity of Rényi divergence across independent mechanisms. This allows for tighter privacy bounds and improved analysis of iterative algorithms like SGD. Compared to traditional DP accounting, RDP can better balance noise addition and utility, making it a valuable tool for scenarios requiring iterative optimization. However, while RDP improves upon the flexibility of privacy composition, its interpretation requires a deeper understanding of Rényi divergence and its relationship to privacy budgets, potentially complicating its application in some domains [73].
2.3 Explainable AI
@@ 3.1 Hyperdimensional Computing
Hyperdimensional Computing (HD) draws significant inspiration from the architecture of the human brain, utilizing high-dimensional vectors to simulate cognitive functions [77, 78, 79]. This advanced approach allows HD to emulate brain-like methodologies for processing, analyzing, and storing information across diverse cognitive tasks. The HD framework, depicted in Figure 1, is structured into four key phases: encoding, training, inference, and retraining. Each phase is meticulously designed to model different aspects of cognitive processing.
@@ 3.3 Differential Privacy-Hyperdimensional Computing
In this work, we present the DP-HD model, a novel approach in AM that integrates DP. This model excels in real-time monitoring and enables for precise mathematical adjustments of noise levels, crucial for safeguarding privacy without compromising accuracy, as illustrated in Algorithm 1. Ensuring the confidentiality of training data in AM is vital to prevent potential security breaches. To address this, Gaussian noise is strategically injected into the class hypervectors, thereby preventing the extraction of individual training samples by unauthorized entities.
During the training phase of the HD model, DP is implemented by making the class hypervectors differentially private through the injection of Gaussian noise, as depicted in Equation (8). This approach effectively reduces the risk of an adversary inferring specific details about the original feature vectors from the class hypervectors. The optimal $\sigma_{dp}$ is determined using Equation (7), which aims to balance data privacy and the operational performance of our predictive model in the AM setting. In our specific case, we set $\delta$ to $10^{-4}$, based on the principle that $\delta$ should be smaller than the inverse of the dataset size. The selection of the appropriate noise level, which is crucial for enhancing the model’s security, depends on the DP-HD model’s sensitivity. This sensitivity is determined by identifying the maximum norm across all encoded training samples, as shown in Equation (9). A higher sensitivity necessitates the incorporation of greater noise levels, thus reinforcing the model’s security. Therefore, the highest sensitivity observed across all training samples is used as a comprehensive sensitivity indicator, ensuring robust privacy measures across the dataset.
@@
Another crucial parameter that influences the privacy level and sets the noise level is the privacy budget. The privacy budget can be selected to predict the impact of noise on the model’s performance. In Section 3.3.1, we delve into the explainability of the DP-HD mechanism. We begin by assessing the effect of noise and training samples on the model’s performance through an examination of its memorization capability. We then define the SNR, which is derived from the ratio of the signal (training samples) contribution to the model’s performance against the noise contribution. This metric enables us to forecast how noise levels affect accuracy, allowing us to set specific privacy and performance levels concurrently. Following this, we explore the decodability of DP-HD. We analyze an analytical method for extracting information from a hypervector by subtracting two secure class hypervectors generated from datasets differing by only one sample. Moreover, in Section 3.3.2, we outline an algorithm to find the optimal encoding strategy to balance privacy and performance. The encoding phase is critical to the model’s success, as it determines how data is mapped from feature space to hyperspace.
@@ 3.3.1 Explainability
Memorization in DP-HD occurs through the bundling operation during the training phase. As described earlier, the memorized vector is given by Equation (1). To ensure DP during training, Gaussian noise with an SD of $\Delta g\sigma_{dp}$ is added to each component of the class hypervectors, as shown in Equation (8). This noise plays a crucial role during the inference process, where the algorithm attempts to recall one of the memorized hypervectors. In the presence of noise, the inference relies primarily on the similarity inner product between the secured class hypervector $\tilde{\vec{C}}_{s}$ and the query hypervector $\vec{H_{q}}$:
@@
This relationship allows us to quantify the impact of noise on the system’s accuracy. A higher SNR indicates better accuracy by emphasizing the signal’s dominance over noise during the inference stage. This equation also shows that increasing $\epsilon$ boosts the SNR, resulting in higher accuracy. However, it also reduces the noise level, thereby decreasing the privacy level according to Equations (7) and (8). Therefore, we select an $\epsilon$ value that balances the required privacy and model performance levels.
Moreover, as we can see, with a very large dimension $D$, the SNR ratio significantly decreases. The only way to counter this effect is to increase $\sigma_{dp}$, which is effectively done through a loss of DP by increasing $\epsilon$ and $\delta$. However, an alternate method of increasing the SNR is by simply increasing the dataset size $N$. Thus, having more training data reduces the effect of DP on the quality of the model. This can be intuitively understood since the data is sampled from the same underlying data distribution, the more dense the training sample is, the less the effect of removing a single training point becomes. In the limit of small ${\rm SNR}$, for good performance the width of the signal distribution should be much smaller than the width of the noise distribution.
@@
$=\mathbb{E}\left((A_{i})^{-1}\vec{\xi}_{i}\vec{\xi}_{i}^{\dagger}(A_{i}^{\dagger})^{-1}\right)$
$=\sqrt{2}\mathbb{E}\left((A_{i})^{-1}\vec{\xi}_{i}\vec{\xi}_{i}^{\dagger}(A_{i}^{\dagger})^{-1}\right)$
@@
$\hat{\vec{F}}=\vec{F}+\mathcal{N}(\vec{0},\Sigma)$ (21)
where $\Sigma=\Sigma^{0}/K={\sqrt{2}\Delta g\sigma_{dp}}/K$. Thus, as expected, the noise during the decoding process also depends linearly on the sensitivity of the encoder and the amount of differential private noise added to the encoder. For a better DP, we need to add more noise which will increase the error in the decoder. This error can be reduced by increasing the value of the sampled in $K$. However, note that since $\Delta g$ is proportional to $D$, and $K$ is proportional to $D$, the ratio $\Delta g/K$ is independent of $D$. Thus, increasing the dimension should not change the error in the decoding process.
@@ 3.3.2 Encoding Strategy
The encoding strategy plays a vital role in balancing privacy and performance in the DP-HD model. To achieve this balance, it is essential to find the optimal SD, denoted as ${\sigma}^{*}_{\text{b}}$, which ensures both the highest accuracy and the highest privacy. The value of ${\sigma}^{*}_{\text{b}}$ is influenced by the distribution of training samples and the distance between the centers of different classes. It is essential to note that ${\sigma}^{*}_{\text{b}}$ is not dependent on the DP-HD model parameters, such as the size of the hypervectors.
To determine ${\sigma}^{*}_{\text{b}}$, we select a subset of training samples that represent the distribution of the entire dataset for training, and another subset for validation. We then experiment with different ${\sigma}_{\text{b}}$ values to generate random basis vectors, as outlined in Algorithm 2. Using these values, we create secured class hypervectors with the training data for various privacy budget values. We then calculate the validation accuracy based on the model’s performance in predicting the validation data labels. We store the accuracy for each pair of (${\sigma}_{\text{b}}$,$\epsilon$). Finally, we select the SD for generating random basis vectors, ${\sigma}^{*}_{\text{b}}$, that corresponds to the highest accuracy and the lowest $\epsilon$.
@@ 3.4.1 Decodability of Hypervectors
Figure 3: The experimental setup, showing (a) the EOS M270 laser powder bed fusion additive manufacturing machine [81], (b) the top view schematic of the built part, and (c) the side view schematic with dimensions, including a 40.5-degree overhang.
@@ 4 Experimental Design
Addressing privacy concerns in the dissemination and analysis of sensor data from AM processes is a paramount consideration. In industrial settings, protecting sensitive manufacturing data, such as part designs and proprietary process parameters, is crucial to maintaining competitive advantage and ensuring regulatory compliance. This study, therefore, not only focuses on the technical aspects of in-situ sensing and defect detection but also realizes the implementation of DP measures. These measures ensure that the valuable insights gained from high-speed camera data do not compromise the confidentiality of the manufacturing process or the design specifications of the parts being produced. This introduction sets the stage for a comprehensive investigation into the experimental designs and methodologies employed in this research. The subsequent sections delve into the specifics of sensor integration, data acquisition, and the analytical techniques used to achieve the dual objectives of enhancing manufacturing precision and ensuring data privacy in the realm of AM.
@@ 4.1 LPBF AM Machine Instrumentation and Setup
This section outlines the instrumentation of a commercial LPBF machine, specifically the EOS M270 model situated at the National Institute of Standards and Technology (NIST) as shown in Figure 3(a). Integration involved outfitting the machine with a high-speed visible camera. The camera boasts a frame rate of 400 frames per second and a spectral response spanning 300-950 nm, facilitating real-time monitoring of the LPBF process. This monitoring entails capturing disparities in the thermal signature of the part as it undergoes layer-wise construction, enhancing quality assurance. Positioned within the build chamber’s upper right corner, the high-speed visible camera is connected via data cables routed through a custom sealed through-port. Notably, precise measurements regarding the camera’s orientation and distance were not conducted. The camera utilized features a Silicon-based array and boasts a resolution of 1.2 megapixels. Its placement inside the build chamber enables comprehensive imaging of the part’s shape and surrounding spatter pattern. Images captured by the high-speed visible camera are windowed to dimensions of 256 pixels × 256 pixels and acquired at a rate of 1000 frames per second. These images serve as representative snapshots, depicting both overhang and bulk build features. The experimental setup replicates conditions in industrial LPBF processes, providing insights into defect detection and quality control practices [82].
@@
The example part was fabricated utilizing specific build parameters tailored for a commercial LPBF system. These parameters include a hatch distance of 0.1 mm, a stripe width of 4 mm, a stripe overlap of 0.1 mm, a powder layer thickness of 20 $\mu$m, a laser scan speed of 800 mm/s, and a laser power of 195 W during infill. Contour passes commence from distinct corners of the part, as indicated in Figure 3(b), traversing counterclockwise along the part perimeter. Laser power settings for these passes vary, with pre-contour passes utilizing 100 W and post-contour passes employing 120 W, both at a scan speed of 800 mm/s. These parameters mirror those used in real-world industrial applications, emphasizing the practical relevance of the study’s findings. The study’s focus lies on acquiring and initially processing sensor data, which entails exploring a single layer example from the 16 mm tall part. Data collection spans 45 different build layers, with the overhang delineated as the last two scan vectors prior to or just after edge formation, excluding pre- or post-contour scans. The remaining scans are attributed to the bulk volume of the part. A stripe pattern scan strategy is adopted (Figure 3(b) and 3(c)), with laser scans over the overhang occurring four times per layer beyond 4 mm build height. The stripe orientation shifts by 90° between layers, and three example layers exhibit a vertical stripe pattern. The fabricated part and its features were chosen to reflect common challenges in industrial AM processes, such as overhang formation and thermal distortions. The test artifact, composed of nickel alloy 625, features a 40.5-degree overhang without additional support structures. Sensor data analysis focuses on three example build heights: 6.06 mm, 7.90 mm, and 9.70 mm, encompassing the formation of the overhang structure. Each layer consists of four 4 mm stripes alternated by 90°. The nickel alloy 625 powder originates from the machine vendor, with a mean powder diameter of 37.8 $\mu$m, utilized in conjunction with a custom alloy 625 substrate. A side-view schematic of the LPBF part (Figure 3(c)) illustrates the investigation into the effects of an overhanging structure. Data exploration in this paper centers on a single build layer at a height of 7.90 mm, although the overhanging structure extends to complete a 16 mm tall object.